OpenStudy (anonymous):

cube root of 108 - cube root of 32

OpenStudy (anonymous):

\[\sqrt[3]{108}-\sqrt[3]{32}=\sqrt[3]{27\times4}-\sqrt[3]{8\times4}\] \[=3\sqrt[3]{4}-2\sqrt[3]{4} =\sqrt[3]{4}\]